# FreeStyle: Free Lunch for Text-guided Style Transfer using Diffusion Models

### v1

@@ Abstract
The rapid development of generative diffusion models has significantly advanced the field of style transfer. However, most current style transfer methods based on diffusion models typically involve a slow iterative optimization process, e.g., model fine-tuning and textual inversion of style concept. In this paper, we introduce FreeStyle, an innovative style transfer method built upon a pre-trained large diffusion model, requiring no further optimization. Besides, our method enables style transfer only through a text description of the desired style, eliminating the necessity of style images. Specifically, we propose a dual-stream encoder and single-stream decoder architecture, replacing the conventional U-Net in diffusion models. In the dual-stream encoder, two distinct branches take the content image and style text prompt as inputs, achieving content and style decoupling. In the decoder, we further modulate features from the dual streams based on a given content image and the corresponding style text prompt for precise style transfer. Our experimental results demonstrate high-quality synthesis and fidelity of our method across various content images and style text prompts. The code and more results are available at our project website: https://freestylefreelunch.github.io/.
Figure 1: Illustration depicting process distinctions among Finetune-based, Inversion-based, and our FreeStyle. (a) Fine-tuning the entire model or specific parameters allows embedding a given visual style into the output domain of the generated diffusion model. (b) Embed specific style or content into a new pseudo-word (e.g., ’[C]’) via training set inversion, and use prompts with this pseudo-word to effectively achieve style transfer. (c) In contrast to the above two methods, FreeStyle requires no optimization and utilizes the intrinsic style reconstruction ability of the diffusion model for effective style transfer.
@@ 1 Introduction
Image style transfer Jing et al. (2019); Gatys et al. (2015) intends to transfer the natural image into the desired artistic image while preserving the content information. With the recent rapid development of generative diffusion models Rombach et al. (2022); Podell et al. (2023), image style transfer has also witnessed significant advancements. These methods can be broadly classified into two categories: finetuning-based methods Wang et al. (2023); Huang et al. (2022) and inversion-based methods Zhang et al. (2023b); Mokady et al. (2023). The former (depicted in Figure 1 (a)) requires optimizing some or all parameters to degrade the model to generate images of specific styles, while the latter (illustrated in Figure 1 (b)) involves learning the specific style concept as the textual token to guide style-specific generation. Both approaches often require thousands or even more iterations of training, leading to significant computational costs and a slow optimization process.
@@
Large text-guided diffusion models Rombach et al. (2022); Zhang et al. (2023a); Saharia et al. (2022), on the other hand, are typically trained on large-scale datasets of text-image pairs, e.g., LAION dataset Schuhmann et al. (2021), which encompasses various style images and corresponding style text prompts. Consequently, these models Rombach et al. (2022); Podell et al. (2023) inherently possess the generative ability for specific styles. This raises the question: How to leverage the various style generation capabilities of pre-trained text-guided diffusion models for style transfer tasks without additional optimization?
@@
Recent works Alaluf et al. (2023); Hertz et al. (2023) have introduced a cross-image attention mechanism to pre-trained diffusion models, enabling the control of appearance or style transfer without any optimization. However, the use of appearance images or style images as references is still required. In some practical applications, users may not have access to reference images, yet they desire to engage in image transfer based on style text prompts. For instance, users can envision transforming their photos into styles reminiscent of Picasso or Da Vinci without possessing works by these renowned artists. In this paper, we present a novel style transfer approach that requires neither optimization nor style images. Specifically, we propose a novel structure composed of a dual-stream encoder and a single-stream decoder. In this configuration, the dual-stream encoder separately encodes the content image and style text prompt as inputs, extracting features from the corresponding modalities for integration in the decoder. We believe images consist of low-frequency signals controlling image content and high-frequency signals governing image style Seo (2020); Shang et al. (2023). Inspired by FreeU Si et al. (2023), which instantiates two modulation factors to balance low-frequency features from the U-Net’s main backbone and high-frequency features from skip connections to improve the quality of image generation. We modulate the feature maps generated by two distinct encoders using two scaling factors, the first scaling factor regulates the strength of style transfer in the image and the second scaling factor controls the degree of content preservation in the images. Our approach is extremely simple and efficient, requiring only the adjustment of appropriate scaling factors to achieve the transfer of a specific style for any image.
[…]
@@
  • We conducted comprehensive experiments with a wide range of images from various artists and styles. The results show that the art images generated by FreeStyle exhibit accurate style expression and high-quality content-style fusion. Compared to state-of-the-art methods, FreeStyle demonstrates superior and more natural stylization effects.
@@ 2.1 Image Style Transfer
The field of style transfer plays a pivotal role in image processing and computer vision. It has seen a rapid evolution from manual texture synthesis Wang et al. (2004); Zhang et al. (2013) to advanced neural style transfer (NST) Jing et al. (2019); Zhang et al. (2019); Sanakoyeu et al. (2018); Deng et al. (2022); Zhang et al. (2022); He et al. (2023), marking a significant shift from traditional techniques to modern deep learning approaches. Generative Adversarial Networks (GANs) Goodfellow et al. (2014), with impressive image generation capabilities, have been rapidly applied to style transfer tasks Zhu et al. (2017); Karras et al. (2019); Gal et al. (2022), further advancing the development of the field. With the recent rapid development of generative diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020), significant progress has been made in image style transfer. These techniques can be classified into two main categories: finetune-based methods and inversion-based methods. Finetune-based methods Wang et al. (2023); Huang et al. (2022) optimize some or all of the model parameters using extensive style images, embedding their visual style into the model’s output domain. In contrast, inversion-based methods Zhang et al. (2023b); Mokady et al. (2023) embed style or content concepts into special word embeddings using style or content images and achieve style transfer with prompts containing these word embeddings. The aforementioned methods based on diffusion models require style images for training models, leading to a slow optimization process. Recent works Alaluf et al. (2023); Hertz et al. (2023) introduce a cross-image attention mechanism and develop a style transfer method that requires any optimization. However, these methods still rely on style images as references. As a text-guided style transfer method, FreeStyle differs by modulating the latent space features of the diffusion model, leveraging its inherent decoupling ability for style transformation without the need for extra optimization or style reference images.
2.2 Text-guided Synthesis
@@
GAN-CLS Reed et al. (2016) is the first to achieve text-guided image synthesis of flowers and birds using recurrent neural networks Lipton et al. (2015) and Generative Adversarial Networks Goodfellow et al. (2014). Subsequently, numerous efforts in text-guided image generation Zhang et al. (2017); Xu et al. (2018); Zhu et al. (2019) have propelled rapid development in this field. Benefiting from the introduction of CLIP Radford et al. (2021), GLIDE Nichol et al. (2021) quickly became the first to implement text-guided image generation that conforms to descriptions, following the development of diffusion models. In addition to generating images that match text descriptions, text-guided techniques are now widely used in various tasks such as image editing Wu et al. (2023b); Gal et al. ; Kawar et al. (2023), image restoration Qi et al. (2023); Lin et al. (2023), and video synthesis Esser et al. (2023) etc. Tsu-Jui Fu et al. Fu et al. (2022) argue that traditional style transfer methods, which depend on pre-prepared specific style images, are both inconvenient and creativity-limiting in practical applications. Following this, a new style transfer method that is guided by textual descriptions Kwon and Ye (2022); Patashnik et al. (2021) has been introduced, offering enhanced flexibility and convenience. This not only simplifies complex Artistic creation but also makes advanced image manipulation accessible to a broader audience without the need for specialized graphic design skills. As a result, text-guided image processing is revolutionizing the way we interact with and create visual content.
2.3 Diffusion Models
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) have recently witnessed remarkable advancements, demonstrating their superiority in image generation Nichol et al. (2021); Ramesh et al. (2022); Saharia et al. (2022), video generation Ho et al. (2022); Wu et al. (2023a), 3D generation Watson et al. (2022); Li et al. (2022), etc. Given the impressive generative capabilities of diffusion models, a considerable amount of style transfer works based on diffusion models Zhang et al. (2023b); Wang et al. (2023); Yang et al. (2023); Wu et al. (2023b); Mokady et al. (2023) have been proposed. To further enhance the quality of image synthesis, SDXL Podell et al. (2023) is trained on data with various aspect ratios, utilizing additional attention modules and a larger cross-attention context. Furthermore, FreeU Si et al. (2023) proposed a method to enhance the quality and fidelity of image generation by leveraging the latent capabilities of the U-Net architecture. Similarly, this work utilizes SDXL Podell et al. (2023) as the backbone network to maximize high-quality image synthesis.
@@ 3.1 Preliminaries
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) involve a forward diffusion process and a reverse denoising process, designed for modeling the mapping between the actual data distribution and the distribution of Gaussian noise. During the forward process, Gaussian noise $\epsilon$ is progressively added to the clean sample $x_{0}$, with the intensity of the added noise $\epsilon$ increasing as $t\in\left[1,2,\dots,T\right]$ increases. The noised image at step $t$ is obtained through the diffusion process:
@@
where $\epsilon\sim\mathcal{N}\left(0,\mathcal{I}\right)$, $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha_{t}}=\prod\limits_{i=0}^{t}\alpha_{i}$, and $\beta_{t}\in\left(0,1\right)$ is a fixed variance schedule. Conversely, in the denoising process, $x_{T}$ is gradually transformed into the clean image $x_{0}$ by progressively predicting and removing the noise. In DDIM Song et al. (2020), to preserve content information, the deterministic sampling from step $t$ to step $t-1$ can be represented as:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}\left(x_{t}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}\left(x_{t},t,c\right).$ (2)
Here, $c$ represents the condition input (e.g., text embedding), and $\epsilon_{\theta}$ denotes the noise prediction network. The predicted denoised image, denoted as $\hat{x}_{0,t}\left(x_{t}\right)$, is obtained by:
$\hat{x}_{0,t}\left(x_{t}\right):=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(x_{t},t,c\right)}{\sqrt{\bar{\alpha}_{t}}}.$ (3)
To reduce computational demands and achieve high-resolution generation, LDM Rombach et al. (2022) applied diffusion denoising operation on the latent space of a pre-trained encoder.
@@ 3.2 Model Structure of FreeStyle
In diffusion models, the U-Net structure is commonly used as the noise prediction network. It consists of an encoder and a decoder, along with skip connections that facilitate information exchange between corresponding layers of the encoder and decoder. Inspired by FreeU Si et al. (2023), which proposes to balance the low-frequency and high-frequency features from the U-Net’s backbone and skip layers, we introduce a novel modulate method of fusing content information and style information applied to style transfer. Figure 2 (a) illustrates the overall structure of FreeStyle, which consists of a dual-stream encoder and a single-stream decoder. The dual-stream encoder in FreeStyle comprises two U-Net encoders with shared parameters, while the single-stream decoder is made up of the U-Net decoder structure. The dual-stream downsampling process can be described separately as follows:
$\begin{cases}f_{s}=&\mathrm{E}\left(x_{\sigma},c\right)\\ f_{c}=&\mathrm{E}\left(x_{0}\right),\\ \end{cases}$ (4)
where $c$ represents the embedding of the style text prompt, and $x_{\sigma}$ denotes the image of the content after $\sigma$ steps of noise addition. $f_{s}$ and $f_{c}$ respectively represent image features that carry style and content information. The denoising process reverses the diffusion process to the predicted clean data $x_{t-1}$ given the nosiy input $x_{t}$:
$p_{\theta}\left(x_{t-1}|x_{t}\right)=$
$\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{0},x_{\sigma},c,t\right),\Sigma_{\theta}\left(x_{0},x_{\sigma},c,t\right)\right).$ (5)
@@
Figure 4: The results of style transfer on portraits using FreeStyle. Under the conditions of fine-grained anime style text prompts, the stylized results achieved with FreeStyle still exhibit clear fine-grained style differences and achieve a natural fusion of style and content.
@@ 3.3 Feature Modulation Module
FreeU Si et al. (2023) strategically reweights the contributions from U-Net’s skip connections and backbone feature maps, effectively harnessing the strengths of these two components of the U-Net architecture to enhance the quality of the generated images. We believe images consist of low-frequency signals controlling image content and high-frequency signals governing image style. Hence, we implement an effective training-free style transfer by modulating the style feature $f_{s}$ and the content feature $f_{c}$. Different from FreeU, two features that need to be modulated draw from two different inputs, style input $f_{s}$ and content input $f_{c}$.
As shown in Figure 2(b), the content feature $f_{c}$ is generated guided by the noise-free content image $x_{0}$, while the style feature $f_{s}$ is generated guided by the style text prompt $c$ and the noise-added image $x_{\sigma}$. During the upsampling process in U-Net, the features $f_{c}$ primarily influence the semantic expression of the generated result, while the features $f_{s}$ have a greater impact on the high-frequency detail information of the result. Consequently, we engage in special modulation of $f_{s}$ and $f_{c}$ to further activate the intrinsic style reconstruction capability of U-Net. To enhance the semantic characteristics of the feature $f_{c}$, we amplify their variance. Specifically, we apply a weight parameter $b$ greater than 1 to certain dimensions of the feature to expand their variance. We can concisely represent this process as follows:
$f_{c}=concat\left(b\times f_{c}\left[:n\right],f_{c}\left[n:\right]\right).$ (6)
In the formula, the $n$ is used to truncate a portion of the features. On the other hand, to extract style features from the feature $f_{s}$, we believe it is necessary to suppress the low-frequency semantic characteristics while preserving high-frequency details and other style expression information. To achieve this, we first transform the feature $f_{s}$ into frequency domain information using the Fourier transform, and then apply a threshold $r_{\text{thresh}}=1$ to filter out the low-frequency semantic information from the features. Subsequently, we use a weight parameter $s$ greater than 1 to enhance the style information. Finally, we convert the processed frequency domain features back into spatial domain features using the inverse Fourier transform. We can simply denote this process as:
$f_{s}=IFFT\left(\mathcal{F}\left(FFT\left(f_{s}\right)\right)\right),$ (7)
@@
where $r$ is the radius. Applying the above method, we modulate the $f_{c}$ and the $f_{s}$, and finally concatenate them to feed into the blocks of U-Net decoder.

### v2

@@ Abstract
The rapid development of generative diffusion models has significantly advanced the field of style transfer. However, most current style transfer methods based on diffusion models typically involve a slow iterative optimization process, e.g., model fine-tuning and textual inversion of style concept. In this paper, we introduce FreeStyle, an innovative style transfer method built upon a pre-trained large diffusion model, requiring no further optimization. Besides, our method enables style transfer only through a text description of the desired style, eliminating the necessity of style images. Specifically, we propose a dual-stream encoder and single-stream decoder architecture, replacing the conventional U-Net in diffusion models. In the dual-stream encoder, two distinct branches take the content image and style text prompt as inputs, achieving content and style decoupling. In the decoder, we further modulate features from the dual streams based on a given content image and the corresponding style text prompt for precise style transfer. Our experimental results demonstrate high-quality synthesis and fidelity of our method across various content images and style text prompts. Compared with state-of-the-art methods that require training, our FreeStyle approach notably reduces the computational burden by thousands of iterations, while achieving comparable or superior performance across multiple evaluation metrics including CLIP Aesthetic Score, CLIP Score, and Preference. We have released the code anonymously at: https://anonymous.4open.science/r/FreeStyleAnonymous-0F9B.
@@ 1 Introduction
Image style transfer [1, 2] intends to transfer the natural image into the desired artistic image while preserving the content information. With the recent rapid development of generative diffusion models [3, 4], image style transfer has also witnessed significant advancements. These methods can be broadly classified into two categories: finetuning-based methods [5, 6] and inversion-based methods [7, 8]. The former (depicted in Figure 1 (a)) requires optimizing some or all parameters to degrade the model to generate images of specific styles, while the latter (illustrated in Figure 1 (b)) involves learning the specific style concept as the textual token to guide style-specific generation. Both approaches often require thousands or even more iterations of training, leading to significant computational costs and a slow optimization process.
@@
Large text-guided diffusion models [3, 9, 10], on the other hand, are typically trained on large-scale datasets of text-image pairs, e.g., LAION dataset [11], which encompasses various style images and corresponding style text prompts. Consequently, these models [3, 4] inherently possess the generative ability for specific styles. Recent works [12, 13] have introduced a cross-image attention mechanism to pre-trained diffusion models, enabling control of appearance or style transfer without optimization. However, the use of appearance images or style images as references is still required. In some applications, users may not have access to reference images but want to engage in image transfer based on style text prompts. For instance, users can envision transforming their photos into styles reminiscent of Picasso or Da Vinci without possessing works by these renowned artists.
@@
In this paper, we present a novel style transfer approach that requires neither optimization nor style images. Specifically, we propose a novel structure composed of a dual-stream encoder and a single-stream decoder. In this configuration, the dual-stream encoder separately encodes the content image and style text prompt as inputs, extracting features from the corresponding modalities for integration in the decoder. It has been demonstrated that the low-frequency signals and high-frequency signals of an image are strongly correlated with its semantic information and style information [14, 15], respectively. We instantiate two modulation factors to balance low-frequency features from the U-Net’s main backbone and high-frequency features from skip connections to implement image style transfer. The first scaling factor regulates the strength of style transfer in the image and the second scaling factor controls the degree of content preservation in the images. Our approach is extremely simple and efficient, requiring only the adjustment of appropriate scaling factors to achieve the transfer of a specific style for any image.
Through strategically modulating feature maps from U-Net’s skip connections and backbone, our FreeStyle framework exhibits seamless adaptability of style transfer when integrated with the existing large text-guided diffusion models, e.g., SDXL [4]. It is important to note that despite structural differences from the U-Net [16] model in pre-trained diffusion models, our approach incorporates U-Net modules without introducing new parameters. To our knowledge, FreeStyle is the first style transfer method based on diffusion models that neither requires reference style images nor any optimization. We conduct a comprehensive comparison of our method with other state-of-the-art techniques, including CLIPstyler [17], CAST [18], StyTr2 [19], UDT2I [20], etc.
Our contributions are summarized as follows:
@@
  • We conduct comprehensive experiments with a wide range of content images and various style texts. The results show that the art images generated by FreeStyle exhibit accurate style expression and high-quality content-style fusion. Compared to state-of-the-art methods, FreeStyle demonstrates superior and more natural stylization effects. In our quantitative experiments, FreeStyle’s CLIP Aesthetic Score improved by 1.4% over others, Preference surpassed others by 32%, and it also showed competitive performance in CLIP Score.
Figure 1: Illustration depicting distinctions among fine-tune-based, inversion-based, and our FreeStyle approaches. (a) Fine-tuning the entire model or specific parameters embeds a visual style into the output domain of the diffusion model. (b) Embedding a specific style or content into a new pseudo-word (e.g., ‘[C]’) via training set inversion, and using prompts with this pseudo-word to achieve style transfer. (c) Unlike the above methods, FreeStyle requires no optimization and utilizes the intrinsic style reconstruction ability of the diffusion model for effective style transfer.
@@ 2 Related Work
Image Style Transfer. The field of style transfer plays a pivotal role in image processing and computer vision. It has seen a rapid evolution from manual texture synthesis [21, 22] to advanced neural style transfer (NST) [1, 23, 24, 25, 19, 18], marking a significant shift from traditional techniques to modern deep learning approaches. Generative Adversarial Networks (GANs) [26], with impressive image generation capabilities, have been rapidly applied to style transfer tasks [27, 28, 29], further advancing the development of the field. With the recent rapid development of generative diffusion models [30, 31], significant progress has been made in image style transfer. These techniques can be classified into two main categories: finetune-based methods and inversion-based methods. Finetune-based methods [5, 6] optimize some or all of the model parameters using extensive style images, embedding their visual style into the model’s output domain. In contrast, inversion-based methods [7, 8] embed style or content concepts into special word embeddings using style or content images and achieve style transfer with prompts containing these word embeddings. The aforementioned methods based on diffusion models require style images for training models, resulting in a slow optimization process. Recent works [12, 13] introduce a cross-image attention mechanism and develop a style transfer method that does not require any optimization. However, these methods still rely on style images as references. As a text-guided style transfer method, FreeStyle differs by modulating features of the diffusion model, leveraging its inherent decoupling ability for style transformation without the need for extra optimization or style reference images.
@@
Text-guided Synthesis. GAN-CLS [32] is the first to achieve text-guided image synthesis of flowers and birds using recurrent neural networks [33] and Generative Adversarial Networks [26]. Subsequently, numerous efforts in text-guided image generation [34, 35, 36] have propelled rapid development in this field. Benefiting from the introduction of CLIP [37], the remarkable generative capabilities of text-to-image models [38, 10] have garnered significant attention from researchers, driven by the advancements in diffusion models. In addition to generating images that match text descriptions, text-guided techniques are now widely used in various tasks such as image editing [20, 39, 40], image restoration [41, 42], and video synthesis [43] etc. Tsu-Jui Fu et al. [44] argue that traditional style transfer methods, which depend on pre-prepared specific style images, are both inconvenient and creativity-limiting in practical applications. Following this, a new style transfer method that is guided by textual descriptions [17, 45] has been introduced, offering enhanced flexibility and convenience. This not only simplifies complex artistic creation but also makes advanced image manipulation accessible to a broader audience without the need for specialized graphic design skills. As a result, text-guided image processing is revolutionizing the way we interact with and create visual content.
Deep Model Fusion. Deep model fusion [46, 47, 48] endeavors to integrate multiple deep neural networks (DNNs) into a singular network, maintaining their inherent capabilities and even surpassing the performance of multi-task training [49, 50]. With the emergence of new large language models (LLMs), such as GPT-3 [51], GPT-4 [52], T5 [53] and BERT [54], there is increasing attention on applying weighted averaging (WA) techniques [55, 56, 57] to these models. For instance, B-tuning [58] utilizes Bayesian learning to calculate posterior prediction distributions, thereby fine-tuning the top-K ranked pre-trained models based on their transferability. Zoo-tuning [59] aggregates the weights of pre-trained models with aligned channels to create a final model adapted to downstream tasks, addressing the high costs associated with migrating large models. For diffusion models, FreeU [60] strategically reweights the contributions of feature maps from U-Net’s skip connections and backbone to effectively enhance the quality of the generated images without any training. Surpassing FreeU, we further analyze and decouple the features of style text and content images. By directly integrating the two encoder branches of pre-trained U-Net in the diffusion model and modulating their features, we enable style image transfer without requiring textual input or additional training.
@@ 3.1 Preliminaries
Diffusion models [30, 31] involve a forward diffusion process and a reverse denoising process. During the forward process, Gaussian noise $\epsilon$ is progressively added to the clean sample $x_{0}$, with the intensity of the added noise $\epsilon$ increasing as $t\in\left[1,2,\dots,T\right]$ increases. The noised image at step $t$ is obtained through the diffusion process:
@@
where $\epsilon\sim\mathcal{N}\left(0,\mathcal{I}\right)$, $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha_{t}}=\prod\limits_{i=0}^{t}\alpha_{i}$, and $\beta_{t}\in\left(0,1\right)$ is a fixed variance schedule. Conversely, in the denoising process, $x_{T}$ is gradually transformed into the clean image $x_{0}$ by progressively predicting and removing the noise. The sampling from step $t$ to step $t-1$ can be represented as:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}\left(x_{t}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}\left(x_{t},t,p\right).$ (2)
Here, $p$ represents the condition input (e.g., text prompt), and $\epsilon_{\theta}$ denotes the noise prediction network.
@@ 3.2 Model Structure of FreeStyle
In diffusion models, the U-Net structure is commonly used as the noise prediction network. It consists of an encoder and a decoder, along with skip connections that facilitate information exchange between corresponding layers of the encoder and decoder. Inspired by FreeU [60], which proposes to balance the low-frequency and high-frequency features from the U-Net’s backbone and skip layers, we introduce a novel modulate method of fusing content information and style information applied to style transfer. Figure 2 (a) illustrates the overall structure of FreeStyle, which consists of a dual-stream encoder and a single-stream decoder. The dual-stream encoder in FreeStyle comprises two U-Net encoders with shared parameters, while the single-stream decoder is made up of the U-Net decoder structure. The dual-stream downsampling process can be described as follows:
$\begin{cases}f_{s}=&\mathrm{E}\left(x_{\sigma},p\right)\\ f_{c}=&\mathrm{E}\left(x_{0}\right),\\ \end{cases}$ (3)
where $p$ represents the embedding of the style text prompt, and $x_{\sigma}$ denotes the content image after $\sigma$ steps of noise addition. The $f_{s}$ and $f_{c}$ represent image features that carry style and content information, respectively. In the denoising process, we predict the noise distribution at step $t$ using the following formula:
$\epsilon_{t}\sim\mathcal{N}\left(\mu_{\theta^{\prime}}\left(f_{c},f_{s},t\right),\Sigma_{\theta^{\prime}}\left(f_{c},f_{s},t\right)\right),$ (4)
where $\theta^{\prime}$ represents the parameters of the decoder in the U-Net, $\mu_{\theta^{\prime}}$ denotes the mean of the predicted noise distribution, and $\Sigma_{\theta^{\prime}}$ indicates the variance of the distribution. Subsequently, we obtain the denoised image $x_{t-1}$ as follows:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}\left(x_{t}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t},$ (5)
where $\hat{x}_{0,t}\left(x_{t}\right)$ represents the estimate of $x_{0}$ given $x_{t}$ and $t$.
@@ 3.3 Feature Modulation Module
We strategically reweight the contributions of feature maps from the encoders of two parameter-sharing U-Nets, effectively leveraging the strengths of both components to implement image style transfer. It has been demonstrated that images consist of low-frequency signals controlling image content and high-frequency signals governing image style [14, 15]. We implement an effective training-free style transfer by modulating the style feature $f_{s}$ and the content feature $f_{c}$ to complete artistic image generation.
As shown in Figure 2(b), the content feature $f_{c}$ is generated guided by the noise-free content image $x_{0}$, while the style feature $f_{s}$ is generated guided by the style text prompt $p$ and the noise-added image $x_{\sigma}$. During the upsampling process in U-Net, the features $f_{c}$ primarily influence the semantic expression of the generated result, while the features $f_{s}$ have a greater impact on the high-frequency detail information of the result. Consequently, we engage in special modulation of $f_{s}$ and $f_{c}$ to further activate the intrinsic style reconstruction capability of U-Net. To enhance the semantic characteristics of the feature $f_{c}$, we amplify their variance. Specifically, we apply a weight parameter $b$ (where $b$>1) to certain channels of the feature to expand their variance, the process can be represent as:
$f_{c}^{\prime}=concat\left(b\times f_{c}\left[:n\right],f_{c}\left[n:\right]\right),$ (6)
where $n$ is the number of truncated channels of the feature, and $f_{c}^{\prime}$ is the enhanced feature. To suppress the low-frequency semantic characteristics while preserving high-frequency details and other style expression information, we first transform the feature $f_{s}$ into frequency domain information using the Fourier transform, and then apply a threshold $r_{\text{thresh}}=1$ to filter out the low-frequency semantic information from the features. Subsequently, we use a weight parameter $s$ greater than 1 to enhance the style information. Finally, we convert the processed frequency domain features back into spatial domain features using the inverse Fourier transform. The process can be simply denoted as:
$f_{s}^{\prime}=IFFT\left(\mathcal{F}\left(FFT\left(f_{s}\right)\right)\right),$ (7)
@@
where $r$ is the radius. By applying the above operations, we modulate $f_{c}$ and $f_{s}$ to obtain $f_{c}^{\prime}$ and $f_{s}^{\prime}$, and finally concatenate $f_{c}^{\prime}$ and $f_{s}^{\prime}$ to feed them into the blocks of the U-Net decoder.